OpenStudy (anonymous):

What are the roots of this equation :( Z'' means (d^2Z/dX^2)) Z''''+Z''+1=0

OpenStudy (mathmale):

This is a differential equation. We don't "find the roots of a differential equation." Instead, we write an "auxiliary equation" and find the roots of that. Such an aux. eqn. would be m^4-m^2+1=0. I'd choose some dummy variable, such as p, and let p=m^2. Then our aux. eqn. would become p^2+p+1=0. Find the roots of that (they are complex). I've found that \[p=\frac{ -1 plusorminus \sqrt{-3} }{ 2 }=m ^{2}.\] Unfortunately, that's going to be difficult to solve for m! I'm afraid i'm going to have to defer to someone else in solving this problem. Good luck!